suppose you commit a crime. the way the judge is going to set your sentence is by pulling a number out of a hat. you know that there are ten 1 year sentences, five 2 year sentences and one 20 year sentences in the hat. what is the probability (in percentage) that you willspend 20 years in prison
Answer #1

It is given that

Number of 1 year sentences =10

Number of 2 year sentences=5

Number of 20 year sentences= 1

.°. total numbers in the hat = 10+5+1

= 16

We have to find the probability of getting 20 year sentence

= P(20 year sentence)

= No. of 20 year sentences/Total no. of sentence

= 1/16

% Probability= (1/16)*100

= 6.25%